Grade 12Algebra

101c0+101c2+101c3....101c50 help me find the sum in binomial expansion

Prasi
9 Years agoGrade 12

1 Answer

Sai Pranav
9 Years ago
We know that
101C0+101C1+101C2+101C3+........+101C50+101C51+101C52+........+101C100+101C101=2101
But 
nCr=nCn-r
Therefore
101C51=101C50
101C52=101C49
.….............
101C100=101C1
 
101C101=101C0
Hence
101C0+101C1+101C2+101C3+........+101C50+101C50+.......+101C1+101C0=2101
Which is
2(101C0+101C1+101C2+101C3+........+101C50)=2101
101C0+101C1+101C2+101C3+........+101C50=\frac{2^{101}}{2}
THUS
101C0+101C1+101C2+101C3+........+101C50=2100